lahh7seven:

Megan’s TV had a base price of $189 and the sales tax was 6.88%. Over the next four years, the TV consumed an average of $0.06 of electricity every day. It also needed repairs twice, costing $29 each time. After four years, Megan got a new TV. What was the lifetime cost of Megan’s TV? (Remember that one in every four years is a leap year.) a. $357.68 b. $318.66 c. $347.66 d. $351.65

harliii:

189 x .0688 = 12.8 + 189 = 201.8 (4 x 360) x .06 = 86.4 (29 x 2) + 86.4 = 144.4 201.8 + 144.4 = x that how u solve it